# Conformalized Bayesian Inference, with Applications to Random Partition Models

### v1

@@ Appendix B Additional experiment 1: unimodal posterior distribution over partitions
When constructing the conformal credible region with 90% coverage, we find that it includes the “true partition” when the two leftmost clusters are merged, but no longer the original three-cluster partition. The credible set nevertheless passes the size diagnostic based on 1,000 randomly generated partitions whose number of clusters follows the posterior distribution of $K$: none of these random partitions fall inside the credible region. Finally, the multimodality analysis in Figure 9 confirms that, under this posterior—corresponding to a prior favoring fewer clusters—the true three-cluster configuration ceases to be a posterior mode, leaving the point estimate as the only identified mode.
@@
Appendix C Additional experiment 2: posterior distribution over mixing measures
@@
where $N(\cdot\mid\mu,\sigma^{2})$ denotes a Gaussian density with mean $\mu$ and variance $\sigma^{2}$, and $\tilde{P}$ is a random probability measure on the space of means and variances endowed with a PY process prior (Perman et al.,, 1992; Pitman and Yor,, 1997). Both the prior and posterior distributions of $\tilde{P}$ place their mass on discrete measures with countably many atoms. When appropriate conditional sampling methods are employed (Papaspiliopoulos and Roberts,, 2008; Walker,, 2007; Kalli et al.,, 2011), posterior samples of $\tilde{P}$ can be obtained directly.
Since the Gibbs sampler used in the clustering experiment of the main text marginalizes out $\tilde{P}$ via latent cluster assignments, we refit the model using the conditional sampler implemented in numpyro (Phan et al.,, 2019; Bingham et al.,, 2019), which employs the NUTS algorithm (Hoffman et al.,, 2014). We apply the CBI pipeline to 5,000 training and 1,000 calibration samples from the posterior distribution of $\tilde{P}$, truncated to 10 atoms. Distances between mixing measures are computed using the Wasserstein-1 distance (Nguyen,, 2013), implemented in the pot library (Flamary et al.,, 2021, 2024).
@@
Figures 10(a) and 10(b) show that the posterior concentrates around a single mixing measure, which is in line with the empirical distribution of the observed velocities. The 90% conformal credible region contains the empirical $k$-means mixing measures with $k=4,5,6$, but excludes that with k=3k=3.999The region also passes a basic size check performed by verifying whether 1,000 mixing measures randomly generated according to the prior are included in the region—none of them is. Compared with the analysis in the main text, where a three-cluster configuration was the highest-KDE point estimate under the VI distance on partitions, this result suggests a subtle distinction: while three clusters best summarize the posterior over partitions, more than three nonzero components are required for the corresponding mixing measure, and hence the implied density, to be plausible under the posterior. This provides a refined perspective on the posterior structure captured by the fitted Bayesian model.
Appendix D Additional experiment 3: posterior distribution over covariance matrices
Our final additional experiment applies CBI to covariance estimation. We consider the four-dimensional Iris dataset (Anderson,, 1936), restricted to the Setosa species. After standardizing the observations $X_{1},\ldots,X_{n}$, we fit a simple Gaussian model of the form
$X_{i}\mid\Sigma\overset{\textnormal{iid}}{\sim}N(0,\Sigma),$
with a conjugate Wishart prior on $\Sigma^{-1}$. We draw 5,000 training and 1,000 calibration posterior samples for $\Sigma$ using numpyro, and apply CBI by measuring distances between covariance matrices according to the operator norm
@@ Appendix D Additional experiment 3: posterior distribution over covariance matrices
Figures 11(a) and 11(b) show that the posterior distribution is concentrated around a single, well-defined mode corresponding to the $\|\cdot\|_{\mathrm{op}}$-KDE point estimate. When constructing the 90% conformal credible region, we find that it contains the empirical covariance matrix, but, encouragingly, none of 1,000 randomly generated covariance matrices drawn from the prior.

### v2

@@ 8 Additional experiment 1: unimodal posterior distribution over partitions
When constructing the conformal credible region with 90% coverage, we find that it includes the “true partition” when the two leftmost clusters are merged, but no longer the original three-cluster partition. The credible set nevertheless passes the size diagnostic based on 1,000 randomly generated partitions whose number of clusters follows the posterior distribution of $K$: none of these random partitions fall inside the credible region. Finally, the multimodality analysis in Figure 9 confirms that, under this posterior—corresponding to a prior favoring fewer clusters—the true three-cluster configuration ceases to be a posterior mode, leaving the point estimate as the only identified representative.
@@
9 Additional experiment 2: Colorectal cancer spatial transcriptomics data (dependent posterior distributions over partitions)
In a recent study, duan2025spatially analyze spatial transcriptomics data on colorectal cancer (lee2020lineage), with the aim of discovering spatially aligned subpopulations of tumor, immune, and stromal cells. Spatial alignment refers to the spatial co-localization of clusters across cell types, where spatial coordinates correspond to the two-dimensional positions of cells within a medical image. To capture this structure, the authors propose a spatially-aligned random partition model that combines a PY Gaussian mixture model for gene expression features with a prior dependence structure that induces spatial co-localization in the latent clustering across cell types. Their MCMC output consists of sampled partitions for each cell type: cells of different types are clustered separately, but the partitions are modeled as dependent through prior-induced spatial alignment. The results of their analysis indeed reveal subpopulations (clusters) of cells from different types that are spatially co-localized.
The left-hand side plots in Figure 11 show the $\mathcal{D}_{VI}$-KDE point estimates for tumor and immune cell partitions. Recall that the model of duan2025spatially performs density estimation and clustering on high-dimensional gene expression features while using spatial coordinates to induce alignment. For visualization purposes, we display partitions only in the spatial domain, since spatial alignment is the focus of the analysis. Both point estimates reveal spatially coherent clusters, in line with the findings of duan2025spatially.
Finally, we use the conformal credible regions (constructed separately for both marginal posteriors) to test a hypothesis of global spatial alignment between the partitions of different cell types. Specifically, given the point estimate for type A cells (either tumor or immune), we spatially translate it to type B cells by matching each type B cell to its closest (spatially) type A cell, and assigning that type B cell to the cluster of the matched type A cell. The right-hand side plots in Figure 11 illustrate these translated partitions. We then check whether each translated partition of type B cells lies within the 90% conformal credible region for the posterior distribution over partitions of type B cells. We emphasize that this constitutes a test of global alignment, as it requires the entire translated partition to lie within the credible region. This is distinct from the goal of duan2025spatially, who adopt a more local notion of alignment seeking to identify subsets of clusters that are spatially co-located, rather than assuming full alignment across all clusters. As a result of our analysis, both spatially translated partitions fall outside the corresponding credible regions, supporting the hypothesis that only subsets of tumor and immune cell clusters are spatially co-located. This is consistent with the findings of duan2025spatially.
10 Additional experiment 3: posterior distribution over mixing measures
@@
where $N(\cdot\mid\mu,\sigma^{2})$ denotes a Gaussian density with mean $\mu$ and variance $\sigma^{2}$, and $\tilde{P}$ is a random probability measure on the space of means and variances endowed with a PY process prior (perman1992size, pitman1997two). Both the prior and posterior distributions of $\tilde{P}$ place their mass on discrete measures with countably many atoms. When appropriate conditional sampling methods are employed (papaspiliopoulos2008retrospective, walker2007sampling, kalli2011slice), posterior samples of $\tilde{P}$ can be obtained directly.
Since the Gibbs sampler used in the clustering experiment of the main text marginalizes out $\tilde{P}$ via latent cluster assignments, we refit the model using the conditional sampler implemented in numpyro (phan2019composable, bingham2019pyro), which employs the NUTS algorithm (hoffman2014nuts). We apply the CBI pipeline to 5,000 training and 1,000 calibration samples from the posterior distribution of $\tilde{P}$, truncated to 10 atoms. Distances between mixing measures are computed using the Wasserstein-1 distance (villani2008optimal, nguyen2013convergence), implemented in the pot library (flamary2021pot, flamary2024pot).
@@
Figures 12(a) and 12(b) show that the posterior concentrates around a single mixing measure, which is in line with the empirical distribution of the observed velocities. The 90% conformal credible region contains the empirical $k$-means mixing measures with $k=4,5,6$, but excludes that with k=3k=3.141414The region also passes a basic size check performed by verifying whether 1,000 mixing measures randomly generated according to the prior are included in the region—none of them is. Compared with the analysis in the main text, where a three-cluster configuration was the highest-KDE point estimate under the VI distance on partitions, this result suggests a subtle distinction: while three clusters best summarize the posterior over partitions, more than three nonzero components are required for the corresponding mixing measure, and hence the implied density, to be plausible under the posterior. This provides a refined understanding of the posterior structure captured by the fitted Bayesian model.
11 Additional experiment 4: posterior distribution over covariance matrices
Our final additional experiment applies CBI to covariance estimation. We consider the four-dimensional Iris dataset (anderson1936species), restricted to the Setosa species. After standardizing the observations $X_{1},\ldots,X_{n}$, we fit a simple Gaussian model of the form
$X_{i}\mid\Sigma\overset{\textnormal{iid}}{\sim}N(\cdot\mid 0,\Sigma),$
with an LKJ prior on the correlation matrix (lewandowski2009generating). We draw 5,000 training and 1,000 calibration posterior samples for $\Sigma$ using numpyro, and apply CBI by measuring distances between covariance matrices according to the operator norm
@@ 11 Additional experiment 4: posterior distribution over covariance matrices
Figures 13(a) and 13(b) show that the posterior distribution is concentrated around a single, well-defined mode corresponding to the $\|\cdot\|_{\mathrm{op}}$-KDE point estimate. When constructing the 90% conformal credible region, we find that it contains the empirical covariance matrix, but, encouragingly, none of 1,000 randomly generated covariance matrices drawn from the prior.